OpenStudy (anonymous):

We know d[x^2]/dx is 2x. However, if we write x^2 as x xs, we get: d[x^2]/dx = d[x+x+ . . .+x (x times)]/dx, i.e. d[x^2]/dx = dx/dx+dx/dx+ . . . +dx/dx (x times), i.e. d[x^2]/dx = 1+1+ . . . +1 (x times), i.e. d[x^2]/dx = x. Where is the logic wrong here?

OpenStudy (anonymous):

product rule d/dx(x.x)=1x+1x=2x

OpenStudy (anonymous):

d/dx of x^3=3x^2 or d/dx (x.x.x) = 1.x.x +x.1.x+x.x.1=3x^2

OpenStudy (anonymous):

I see, makes sense. Thanks